In a hydraulic press, the areas of small and the large pistons are 4cm2 and 600cm2 respectively. Find the velocity ratio of the machine. Assuming there is no friction, find the mechanical advantage.


1
Expert's answer
2021-12-27T15:49:36-0500

Answer

Given data as below

Area of small piston

A =4 cm2=0.0004m2

Area of large post on

A'=600cm2=0.06m2

Now ratio of velocities

Using continuity equation

"Av=A'v'\\ =>\frac{v}{v'}=\frac{A'}{A}"

"=>\frac{v}{v'}=\frac{0.06}{0.0004}"

"=>\frac{v}{v'}=150"


Now maechnical advantage is ratio of output power and input power of mechnical system.